Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4c)
tomelleri - 23:04 Monday 13 October 2025 (67926) Print this report
Operator Report - Afternoon shift

ITF found locked in LOW_NOISE_3 with SCIENCE mode (Autoscience ON); BNS Range ~55 Mpc. Planned commissioning activity for today (times in UTC):
15:37 - ITF in COMMISSIONING mode for scheduled tests on ARM actuators and ETM marionettas. 16:01 - Unlock during WI MAR tests. 16:04 - ITF set to DOWN, continuing tests with ITF unlocked. 16:41 - Restored BS_MAR_TX (+1 urad) to last lock value in order to reach LOCKED_ARMS_IR, parked for a few minutes before continuing lock acquisition at 16:48. 17:04 - Turned off lights in W tunnel. 17:48 - After multiple unlocks each in LOCKING_CARM_NULL_3F,  LOCKING_CARM_NULL_1F and CARM_NULL_1F, ITF request set to ACQUIRE_DRMI to check alignment of SR (-1.5 urad TX) as agreed with Bersanetti.
Relock (UTC): 19:17 - ITF in TROUBLESHOOTING mode since it kept unlocking consistently ~30 seconds after reaching CARM_NULL1_F. 19:30 - After Diego adjusted DIFFp_TY gain (entry #67929) we managed to achieve stable lock in CARM_NULL_1F. 19:59 - ITF back to LN3, in SCIENCE mode at 20:02 with Autoscience engaged; BNS Range ~54 Mpc.

Guard tour (UTC)
17:59 -> 18:35

Sub-system reports

DAQ
14h13m43 UTC - VPM>FbTrend100s_100000 reported Unexpected loss of connection.

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×